# Practical Evaluation of Copula-based Survival Metrics: Beyond the Independent Censoring Assumption

### v1

@@ 6 Conclusion
In this paper, we have proposed three copula-based metrics for evaluating survival models under dependent censoring. We empirically compared them to existing, independent variants in both synthetic and semi-synthetic datasets. We found that in datasets with adequate data and censoring, the dependent metrics can give error estimates that are closer to the true error, mainly in terms of predictive accuracy, i.e., Brier Score and MAE. We anticipate that others will be able to use these metrics and improve upon our framework.
@@ Appendix B Notation
$c^{(i)}$ Censoring time of instance $i$
$e^{(i)}$ Event time of instance $i$
$t^{(i)}$ Observed time of instance $i$
$\hat{t}^{(i)}$ Predicted time of instance $i$
$\delta^{(i)}_{k}$ Event bit/indicator, $\delta^{(i)}=1_{e^{(i)}>c^{(i)}}$
$f_{T}$ Probability density function, representing $\text{Pr}(T=t)$
$F_{T}$ Cumulative density function, representing $\text{Pr}(T<t)$
$S_{T}$ Survival function, $S:\mathbb{R}\rightarrow[0,1]$
@@
$\tilde{\mathcal{D}}$ Raw dataset with flipped censor bit
@@
$h(\cdot\mid\bm{x}^{(i)})$ Hazard function given the covariates $\bm{x}^{(i)}$
@@ Appendix C Censoring assumptions
Non-informative censoring. Censoring can be non-informative if the distribution of survival times ($T_{E}$) provides no information about the distribution of censoring times ($T_{C}$), and vice versa. Otherwise, the censoring is informative. However, we would still need to know which subjects are censored or not censored. The assumption of non-informative censoring is often justifiable when censoring is independent and/or random [Kleinbaum and Klein, 2012, Ch. 1], but cases exist where the censoring is informative but also random and independent: Consider a cohort study, where each time a patient experiences an event (e.g., death), another patient in the study is randomly selected to be censored. This could be a family member who decides to leave the study after an event occurs. If those who are censored are representative of those who remain in the risk set, then the censoring would be random and independent. However, in this case, the censoring would also be informative, as the censoring mechanism would be linked to the time-to-event distribution (since events trigger censoring). In fact, if this were the only mechanism by which individuals were censored, the distribution of survival times would entirely determine the distribution of censoring times, making it highly informative.
@@ Appendix D Copulas and Definitions
We can view bivariate survival analysis as a competing risks problem, where the occurrence of an event (e.g., death) and censoring (e.g., dropout) are mutually-exclusive incidents. Tsiatis [1975] called this the identifiability issue; competing risks data do not allow one to observe two event times simultaneously (one event censors the other), and hence, the data may not identify the dependence structure between event times. Zheng and Klein [1995] gave a partial solution to the identifiability problem of dependent censoring by an assumed copula between two event times. A copula is a function that links two or more random variables by specifying their dependence structure [Emura and Chen, 2018, Ch. 1]. A mathematician, Abe Sklar, first used the word copula in his study of probabilistic metric space [Sklar, 1959]. In his paper, he gave a mathematical definition of copulas and established the most fundamental theorem about copulas, known as Sklar’s theorem [Sklar, 1959], and demonstrated that any joint cumulative density can be written in terms of a copula over the quantiles of its marginal cumulative densities. Among the various types of copulas, the Archimedean copula is the most used, and enables the modeling of dependencies in arbitrarily high dimensions using a single parameter. This family of copulas is defined by a convex generator function, $\varphi:[0,1]\to[0,\infty]$, which is continuous, strictly decreasing, and satisfies the boundary conditions $\varphi(0)=\infty$ and $\varphi(1)=0$ [Emura and Chen, 2018, Ch. 3]. Given $T$ and $C$, we can define an Archimedean copula as:
$\begin{split}P(T>t,C>c)&=\varphi^{-1}\big{[}\varphi\{S_{T}(t)\}\\ &\quad+\varphi\{S_{C}(c)\}\big{]},\end{split}$ (D.1)
where $\varphi$ is the generator function, and $S_{T}(t)=P(T>t)$ and $S_{C}(c)=P(C>c)$ are the marginals survival functions. In the bivariate case, $\varphi$ describes the dependency structure (i.e., correlation) between two random variables, e.g.$T_{E}$ and $T_{C}$, and the (rank) correlation can described with Kendall’s tau ($\tau$):
$\begin{split}\tau=&\quad\text{Pr}\{\ (T_{2}-T_{1})(C_{2}-C_{1})\geq 0\}\\ &-\text{Pr}\{\ (T_{2}-T_{1})(C_{2}-C_{1})<0\},\end{split}$ (D.2)
where ($T_{1}$, $C_{1}$) and ($T_{2}$, $C_{2}$) are sampled from a (survival) copula. The Clayton [Clayton, 1978] and Frank [Frank, 1979] copulas are two common Archimedean copulas in survival literature. Within these families, the copula $C_{\theta}$ is parameterized by $\theta$, interpreted as the degree of dependence between $T_{E}$ and $T_{C}$ in the bivariate case. When $\theta\approx 0$, it yields the independence copula, thus a larger value of $\theta$ implies greater dependency between the marginal distributions. Formally, we write $C(u_{1},\ldots,u_{d}):[0,1]^{d}\rightarrow[0,1]$, which is a $d$-dimensional copula (i.e., where $d$ is the number of variables) with $u_{1},\ldots,u_{d}$ uniform marginal probabilities and the following properties Nelsen [2006]:
@@ 2. Uniform Margins: For all $i\in\{1,\dots,m\}$, if $\forall j\neq i\Rightarrow u_{j}=1$, then
Let $T_{E}$ be the event time, $T_{C}$ is the censoring time, $\bm{x}$ is a vector of covariates, and $S_{T_{E}}(t\mid\bm{x})=\text{Pr}(T_{E}>t_{e}\mid\bm{x})$ and $S_{T_{C}}(u\mid\bm{x})=\text{Pr}(T_{C}>t_{c}\mid\bm{x})$ are the marginal survival functions given $\bm{x}$. We can then define a bivariate (survival) copula $C_{\theta}$ [Nelsen, 2006] that describes the degree of dependence between $T_{E}$ and $T_{C}$ as follows:
$\text{Pr}(T_{E}>t_{e},T_{C}>t_{c}\mid\mathbf{x})\ =\ C_{\theta}\{\,S_{T_{E}}(t_{e}\mid\bm{x}),\,S_{T_{C}}(t_{c}\mid\bm{x})\,\}.$ (D.3)
Kendall’s tau ($\tau$) is a well-known measure to assess the dependence between $T_{E}$ and $T_{C}$, which is defined by [Emura and Chen, 2018, Ch. 3]:
$\begin{split}\tau=&\quad\text{Pr}\{\ (T_{2}-T_{1})(C_{2}-C_{1})\geq 0\mid\bm{x}\ \}-\text{Pr}\{\ (T_{2}-T_{1})(C_{2}-C_{1})<0\mid\bm{x}\ \},\end{split}$ (D.4)
where ($T_{1}$, $C_{1}$) and ($T_{2}$, $C_{2}$) are sampled from the model (D.3). Kendall’s $\tau$ can be solely expressed as a function of the copula $C_{\theta}$ by [Emura and Chen, 2018, Ch. 3]:
@@
which implies that Kendall’s $\tau$ does not depend on the marginal survival functions. An Archimedean copula is a type of copula that is defined using a generator function, which allows for a flexible way to model dependence structures between random variables, e.g.$T_{E}$ and $T_{C}$. Such a copula is defined as [Emura and Chen, 2018, Ch. 3]:
@@
    $C_{\theta}(u,v)=-\frac{1}{\theta}\log\left[1+\frac{(e^{-\theta u}-1)(e^{-\theta v}-1)}{e^{-\theta}-1}\right].$
@@ E.2 Integrated Brier Score
where $t_{\text{max}}$ is usually the maximum event time of the combined training and validation datasets. This metric is implemented in package SurvivalEVAL [Qi et al., 2023b].
@@ E.3 Mean Absolute Error
Additionally, for each censored subject, Haider et al. [2020] suggested we use a confidence weight $\omega_{i}=1-S_{\text{KM}}(t_{c,i})$ for the error calculated based on the margin value. This weight $\omega_{i}$ yields lower confidence for early censoring subjects and higher confidence for late censoring data. Of course, we set the weights for uncensored subjects $i$ to $w_{i}=1$ as we have full confidence in those error calculations. The overall MAE margin after a re-weighting scheme is:
$\text{MAE}_{\text{margin}}\left(\{\bm{x}_{i},t_{i},\delta_{i}\}_{i=1}^{N};\,\hat{t}_{i}\right)=\frac{1}{\sum_{i=1}^{N}\omega_{i}}\sum_{i=1}^{N}\omega_{i}\left|[(1-\delta_{i})\cdot t_{e}^{\text{margin}}(t_{i})+\delta_{i}\cdot t_{i}]-\hat{t}_{i}\right|\ .$
This metric is implemented in package SurvivalEVAL [Qi et al., 2023b].
@@ F.1 Data generation and preprocessing
Table 3: Overview of the raw datasets. The datasets are sorted by the number of instances. The number of features before performing one-hot encoding, with brackets (the number of features after one-hot encoding).
@@
Synthetic: This paper utilizes a linear synthetic dataset with two competing events (i.e., event and censoring), both based on the semi-parametric Weibull distribution as discussed in Foomani et al. [2023]. The datasets are characterized by the conditional hazard function and conditional survival probability of an event as follows:
$h_{T\mid X}(t\mid X)$ $=\left(\frac{v}{\rho}\right)\left(\frac{t}{\rho}\right)^{v-1}\exp\left(g_{\Psi}(X)\right),$ (F.1)
$S_{T\mid X}(t\mid X)$ $=\exp\left(-\left(\frac{t}{\rho}\right)^{v-1}\exp\left(g_{\Psi}(X)\right)\right),$ (F.2)
where $v$ and $\rho$ represent the shape and scale parameters of the Weibull distribution, respectively. We use $v_{E}=4$, $v_{C}=6$, $\rho_{E}=17$, $\rho_{C}=19$. $g_{\Psi}(X)$ represents the risk function, which is a linear function in this work. The data generation process begins by initializing a copula with a predefined parameter $\theta$. We then sample $N$ sets of CDF observations from the copula, each containing two CDF probabilities (corresponding to each competing event). From these probabilities, we then calculate the corresponding times using the inverse function of (F.2). The data generation process is described in Algorithm 2 in Foomani et al. [2023].
@@
SEER: The SEER Program dataset [LAG et al., 2007] is a comprehensive collection of cancer patient data from approximately 49% of the U.S. population. Our study focuses on three subsets: SEER-brain, SEER-liver, and SEER-stomach, containing data on patients diagnosed with brain, liver, and stomach cancers, respectively. The goal is to model the time from diagnosis to failure events, such as death or disease progression, after preprocessing steps outlined by [Qi et al., 2024]. Features analyzed include age, sex, cancer stage, grade, and various clinical factors. We excluded features with over 70% missing data, patients without follow-up times, duplicates, and those with a survival time of zero. The SEER cohort is available for download at www.seer.cancer.gov.
@@ F.2 Survival learners
In this section, we will describe the implementation and hyperparameters of the survival learners.
CoxPH: The Cox Proportional Hazards model (CoxPH) is a semi-parametric survival analysis model [Cox, 1972], consisting of a non-parametric baseline hazard function $h_{0}(t)$ and a parametric partial hazard function $f(\bm{\theta},\bm{x}_{i})$, typically estimated using the Breslow method [Breslow, 1975]. The hazard function is $h(t|\bm{x}_{i})=h_{0}(t)\exp(f(\bm{\theta},\bm{x}_{i}))$, where $f$ is linear. The MLE of $\hat{\bm{\theta}}$ maximizes the Cox partial log-likelihood [Cox, 1972].
GBSA: GBSA [Ridgeway, 1999] is a likelihood-based boosting model that optimizes an $\ell_{2}$-norm penalized Cox partial log-likelihood [Cox, 1972]. It updates candidate variables flexibly at each boosting step using an offset-based gradient boosting method, unlike traditional gradient boosting, which updates either a single component or all features for fitting the gradient.
RSF: Random Survival Forests (RSF) [Ishwaran et al., 2008] extend decision trees for survival analysis by constructing multiple trees on bootstrapped samples. Each tree predicts an individual’s survival function, with splits based on criteria like the log-rank test. RSF handles censored data as right-censored and accounts for it during tree construction, with final predictions as the average of all trees’ survival functions.
DeepSurv: DeepSurv [Katzman et al., 2018] is a neural network implementation of the CoxPH model of the form $h\left(t|\bm{x}_{i}\right)=h_{0}\left(t\right)\exp\left(f\left(\bm{\theta},\bm{x}_{i}\right)\right)$, where $f\left(\bm{\theta},\bm{x}_{i}\right)$ denotes a risk score as a nonlinear function of the features, i.e.$f\left(\bm{\theta},\bm{x}_{i}\right)=\sigma(\bm{x}_{i}\bm{\theta})$, where $\sigma$ is a nonlinear activation function. Popular functions include the hyperbolic tangent (Tanh) function or the Rectified Linear Unit (ReLU). The MLE for $\hat{\bm{\theta}}$ is derived by numerically maximizing the partial Cox log-likelihood.
MTLR: Multi-Task Logistic Regression (MTLR) Yu et al. [2011] is a neural network implementation of a discrete-time model that predicts the survival distribution using a sequence of dependent logistic regressors. The number of discrete times is determined by the square root of the number of uncensored patients, and we use quantiles to divide those uncensored instances evenly into each time interval.

### v2

@@ 6 Conclusion
In this paper, we have proposed three copula-based metrics for evaluating survival models under dependent censoring. We empirically compared them to existing, independent variants in both synthetic and semi-synthetic datasets. We found that in datasets with adequate instances and censoring, the dependent metrics gave error estimates that were closer to the true error for prediction accuracy, i.e., Brier Score and MAE. We anticipate that others will be able to use these metrics and improve upon our framework.
@@ Appendix B Notation
$c_{i}$ Censoring time of instance $i$
$e_{i}$ Event time of instance $i$
$t_{i}$ Observed time of instance $i$
$\hat{t}_{i}$ Predicted time of instance $i$
$\delta_{i}$ Event bit/indicator, $\delta_{i}=\mathbbm{1}[e_{i}>c_{i}]$
$f_{E}$ Probability density function, representing $\text{Pr}(E=t)$
$F_{E}$ Cumulative density function, representing $\text{Pr}(E<t)$
$S_{E}$ Survival function, $S:\mathbb{R}\rightarrow[0,1]$
@@
$\tilde{\mathcal{D}}$ Raw dataset with flipped event bit
@@
$h(\cdot\mid\bm{x}_{i})$ Hazard function given the covariates $\bm{x}_{i}$
@@ Appendix C Censoring assumptions
Non-informative censoring. Censoring can be non-informative if the distribution of survival times ($E$) provides no information about the distribution of censoring times ($C$), and vice versa. Otherwise, the censoring is informative. However, we would still need to know which subjects are censored or not censored. The assumption of non-informative censoring is often justifiable when censoring is independent and/or random [Kleinbaum and Klein, 2012, Ch. 1], but cases exist where the censoring is informative but also random and independent: Consider a cohort study, where each time a patient experiences an event (e.g., death), another patient in the study is randomly selected to be censored. This could be a family member who decides to leave the study after an event occurs. If those who are censored are representative of those who remain in the risk set, then the censoring would be random and independent. However, in this case, the censoring would also be informative, as the censoring mechanism would be linked to the time-to-event distribution (since events trigger censoring). In fact, if this were the only mechanism by which individuals were censored, the distribution of survival times would entirely determine the distribution of censoring times, making it highly informative.
@@ Appendix D Copulas and Definitions
We can view bivariate survival analysis as a competing risks problem, where the occurrence of an event (e.g., death) and censoring (e.g., dropout) are mutually-exclusive incidents. Tsiatis [1975] called this the identifiability issue; competing risks data do not allow one to observe two event times simultaneously (one event censors the other), and hence, the data may not identify the dependence structure between event times. Zheng and Klein [1995] gave a partial solution to the identifiability problem of dependent censoring by an assumed copula between two event times. A copula is a function that links two or more random variables by specifying their dependence structure [Emura and Chen, 2018, Ch. 1]. A mathematician, Abe Sklar, first used the word “copula” in his study of probabilistic metric space [Sklar, 1959]. In his paper, he gave a mathematical definition of copulas and established the most fundamental theorem about copulas, known as Sklar’s theorem [Sklar, 1959]. It demonstrated that any joint cumulative density can be written in terms of a copula over the quantiles of its marginal cumulative densities. Among the various types of copulas, the Archimedean copula is the most used, and enables the modeling of dependencies in arbitrarily high dimensions using a single parameter. This family of copulas is defined by a convex generator function, $\varphi:[0,1]\to[0,\infty]$, which is continuous, strictly decreasing, and satisfies the boundary conditions $\varphi(0)=\infty$ and $\varphi(1)=0$ [Emura and Chen, 2018, Ch. 3]. Given event time $E$ and censoring time $C$, we can define an Archimedean copula as:
$\begin{split}P(E>t,C>c)=\varphi^{-1}\big{[}\varphi\{S_{E}(t)\}+\varphi\{S_{C}(t)\}\big{]},\end{split}$ (D.1)
where $\varphi$ is the generator function, and $S_{E}(t)=P(E>t)$ and $S_{C}(t)=P(C>t)$ are the marginal survival functions. In the bivariate case, $\varphi$ describes the dependency structure (i.e., correlation) between two random variables, e.g.$E$ and $C$, and the (rank) correlation can described with Kendall’s tau ($\tau$):
$\begin{split}\tau=&\quad\text{Pr}\{\ (e_{2}-e_{1})(c_{2}-c_{1})\geq 0\}-\text{Pr}\{\ (e_{2}-e_{1})(c_{2}-c_{1})<0\},\end{split}$ (D.2)
where ($e_{1}$, $c_{1}$) and ($e_{2}$, $c_{2}$) are sampled from a (survival) copula. The Clayton [Clayton, 1978] and Frank [Frank, 1979] copulas are two common Archimedean copulas in survival literature. Within these families, the copula $C_{\theta}$ is parameterized by $\theta$, interpreted as the degree of dependence between ${E}$ and ${C}$ in the bivariate case. When $\theta\approx 0$, it yields the independence copula, thus a larger value of $\theta$ implies greater dependency between the marginal distributions. Formally, we write $C(u_{1},\ldots,u_{d}):[0,1]^{d}\rightarrow[0,1]$, which is a $d$-dimensional copula (i.e., where $d$ is the number of variables) with $u_{1},\ldots,u_{d}$ uniform marginal probabilities and the following properties Nelsen [2006]:
@@ 2. Uniform Margins: For all $i\in\{1,\dots,m\}$, if $\forall j\neq i\Rightarrow u_{j}=1$, then
Let ${E}$ be the event distribution, ${C}$ the censoring distribution, $\bm{X}$ a vector of covariates, and $S_{E\mid\bm{X}}(t\mid\bm{X})=\text{Pr}(E>t\mid\bm{X})$ and $S_{C\mid\bm{X}}(t\mid\bm{X})=\text{Pr}(C>t\mid\bm{X})$ the marginal survival functions given $\bm{X}$. We can then define a bivariate (survival) copula $C_{\theta}$ [Nelsen, 2006] that describes the degree of dependence between $E$ and $C$ as follows:
$\text{Pr}(E>t,C>t\mid\bm{X})\ =\ C_{\theta}\{\,S_{E\mid\bm{X}}(t\mid\bm{X}),\,S_{C\mid\bm{X}}(t\mid\bm{X})\,\}.$ (D.3)
Kendall’s tau ($\tau$) can be similarly defined as:
$\begin{split}\tau=&\quad\text{Pr}\{\ (e_{2}-e_{1})(c_{2}-c_{1})\geq 0\mid\bm{X}\ \}-\text{Pr}\{\ (e_{2}-e_{1})(c_{2}-c_{1})<0\mid\bm{X}\ \},\end{split}$ (D.4)
where ($e_{1}$, $c_{1}$) and ($e_{2}$, $c_{2}$) are sampled from the model (D.3). Kendall’s $\tau$ can be solely expressed as a function of the copula $C_{\theta}$ by [Emura and Chen, 2018, Ch. 3]:
@@
which implies that Kendall’s $\tau$ does not depend on the marginal survival functions. An Archimedean copula is a type of copula that is defined using a generator function, which allows for a flexible way to model dependence structures between random variables, e.g.$E$ and $C$. Such a copula is defined as [Emura and Chen, 2018, Ch. 3]:
@@
    $C_{\theta}(u,v)=-\frac{1}{\theta}\log\left[1+\frac{(\exp(-\theta u)-1)(\exp(-\theta v)-1)}{e^{-\theta}-1}\right].$
@@ E.2 Integrated Brier Score
where $t_{\text{max}}$ is usually the maximum event time of the combined training and validation datasets. This metric is implemented in the SurvivalEVAL package [Qi et al., 2023b].
@@ E.3 Mean Absolute Error
Additionally, for each censored subject, Haider et al. [2020] suggested we use a confidence weight $\omega_{i}=1-S_{\text{KM}}(t_{i})$ for the error calculated based on the margin value. This weight $\omega_{i}$ yields lower confidence for early censoring subjects and higher confidence for late censoring data. Of course, we set the weights for uncensored subjects $i$ to $w_{i}=1$ as we have full confidence in those error calculations. The overall MAE margin after a re-weighting scheme is:
$\text{MAE}_{\text{margin}}\left(\{\bm{x}_{i},t_{i},\delta_{i}\}_{i=1}^{N};\,\hat{t}_{i}\right)=\frac{1}{\sum_{i=1}^{N}\omega_{i}}\sum_{i=1}^{N}\omega_{i}\left|[(1-\delta_{i})\cdot e^{\text{margin}}(t_{i})+\delta_{i}\cdot t_{i}]-\hat{t}_{i}\right|\ .$
This metric is implemented in SurvivalEVAL package [Qi et al., 2023b].
@@ F.1 Data generation and preprocessing
Table 3: Overview of the raw datasets. The datasets are sorted by name. The number of features before performing one-hot encoding, with brackets (the number of features after one-hot encoding).
@@
Synthetic: This paper utilizes a linear synthetic dataset with two competing events (i.e., event and censoring), both based on the parametric Weibull distribution as discussed in Foomani et al. [2023]. The datasets are characterized by the conditional hazard function and conditional survival probability of an event or censoring as follows:
$h_{E\mid\bm{X}}(t\mid\bm{X})$ $=\left(\frac{v}{\rho}\right)\left(\frac{t}{\rho}\right)^{v-1}\exp\left(g_{\Psi}(\bm{X})\right),$ (F.1)
$S_{E\mid\bm{X}}(t\mid\bm{X})$ $=\exp\left(-\left(\frac{t}{\rho}\right)^{v-1}\exp\left(g_{\Psi}(\bm{X})\right)\right),$ (F.2)
where $v$ and $\rho$ represent the shape and scale parameters of the Weibull distribution, respectively. We use $v_{E}=4$, $v_{C}=6$, $\rho_{E}=17$, $\rho_{C}=19$. $g_{\Psi}(\bm{X})$ represents the risk function, which is a linear function in this work. The data generation process begins by initializing a copula with a predefined parameter $\theta$. We then sample $N$ sets of CDF observations from the copula, each containing two CDF probabilities (corresponding to each competing event, i.e., event or censoring). From these probabilities, we then calculate the corresponding times using the inverse function of (F.2). The data generation process is further described in Algorithm 2 in Foomani et al. [2023].
@@
SEER: The SEER Program dataset [LAG et al., 2007] is a comprehensive collection of cancer patient data from approximately 49% of the U.S. population. Our study focuses on three subsets: SEER-brain, SEER-liver, and SEER-stomach, containing data on patients diagnosed with brain, liver, and stomach cancers, respectively. The goal is to model the time from diagnosis to failure events, such as death or disease progression. We preprocess the dataset as in [Qi et al., 2024]. Features analyzed include age, sex, cancer stage, grade, and various clinical factors. We excluded features with over 70% missing data, patients without follow-up times, duplicates, and those with a survival time of zero. The SEER cohort is available for download at www.seer.cancer.gov.
@@ F.2 Survival learners
In this section, we describe the implementation and hyperparameters of the survival learners.
CoxPH: The Cox Proportional Hazards model (CoxPH) is a semi-parametric survival analysis model [Cox, 1972], consisting of a non-parametric baseline hazard function $h_{0}(t)$ and a parametric partial hazard function $f(\bm{\theta},\bm{x}_{i})$, typically estimated using the Breslow method [Breslow, 1975]. The hazard function is $h(t\mid\bm{x}_{i})=h_{0}(t)\exp(f(\bm{\theta},\bm{x}_{i}))$, where $f$ is linear. The MLE of $\hat{\bm{\theta}}$ maximizes the Cox partial log-likelihood [Cox, 1972]. CoxPH is implemented from the scikit-survival package version 0.21.0.
GBSA: GBSA [Ridgeway, 1999] is a likelihood-based boosting model that optimizes an $\ell_{2}$-norm penalized Cox partial log-likelihood [Cox, 1972]. It updates candidate variables flexibly at each boosting step using an offset-based gradient boosting method, unlike traditional gradient boosting, which updates either a single component or all features for fitting the gradient. GBSA is implemented from the scikit-survival package version 0.21.0.
RSF: Random Survival Forests (RSF) [Ishwaran et al., 2008] extend decision trees for survival analysis by constructing multiple trees on bootstrapped samples. Each tree predicts an individual’s survival function, with splits based on criteria like the log-rank test. RSF handles censored data as right-censored and accounts for it during tree construction, with final predictions as the average of all trees’ survival functions. RSF is implemented from the scikit-survival package version 0.21.0.
DeepSurv: DeepSurv [Katzman et al., 2018] is a neural network implementation of the CoxPH model of the form $h\left(t\mid\bm{x}_{i}\right)=h_{0}\left(t\right)\exp\left(f\left(\bm{\theta},\bm{x}_{i}\right)\right)$, where $f\left(\bm{\theta},\bm{x}_{i}\right)$ denotes a risk score as a nonlinear function of the features, i.e.$f\left(\bm{\theta},\bm{x}_{i}\right)=\sigma(\bm{x}_{i}\bm{\theta})$, where $\sigma$ is a nonlinear activation function. Popular functions include the hyperbolic tangent (Tanh) function or the Rectified Linear Unit (ReLU). The MLE for $\hat{\bm{\theta}}$ is derived by numerically maximizing the partial Cox log-likelihood. DeepSurv is implemented from https://github.com/shi-ang/BNN-ISD/.
MTLR: Multi-Task Logistic Regression (MTLR) Yu et al. [2011] is a neural network implementation of a discrete-time model that predicts the survival distribution using a sequence of dependent logistic regressors. The number of discrete times is determined by the square root of the number of uncensored patients, and we use quantiles to divide those uncensored instances evenly into each time interval. MTLR is implemented from https://github.com/shi-ang/BNN-ISD/.